Factor completely.
64x^2 - 48xy + 9y^2

Answer by richwmiller(17219)

You might guess that the first term will have an 8 and the last term will have a 3
since 8^2=64 and 3^2=9
But what next ? how do you get the xy?
obviously there must be y in the second term also
and the second term also must be negative
so what do we know so far.
8x is a good start for the first term
-3y looks like it might be the second term
since (-3y)*(-3y)=9y^2
But will it give us the -48xy
(8x-3y)*(8x-3y)=64x^2-24xy-24xy+9y^2
It works since -24xy-24xy=-48xy
So our answer is
(8x-3y)*(8x-3y) or (8x-3y)^2